Using the 2k ≥ n ​rule, determine the number of classes needed for the following data set sizes.

a) n = 75

b) n = 250

c) n = 1250

d) 2500

a) The number of classes needed when n = 75

b) The number of classes needed when n = 250

c) The number of classes needed when n = 1250

d) The number of classes needed when n = 2500
Answer #1

Answer)

Rule says that 2^k should be greater than or equal to the n

A)

N = 75

We know that 2^7 = 128 which is greater than 75

So k = 7

That is number of classes = 7

B)

N = 250

2^8 = 256 > 250

Number of classes = 8

C)

N = 1250

2^11 = 2048 > 1250

Number of classes = 11

D)

N = 2500

2^12 = 4096 > 2500

Number of classes = 12